When an airplane leaves the runway, its angle of climb is 17° and its speed is 345 feet per second. How long will it take the plane to climb to an altitude of 10,000 feet? 16,000 feet?
Answer by stanbon(75887)

When an airplane leaves the runway, its angle of climb is 17° and its speed is 345 feet per second. How long will it take the plane to climb to an altitude of 10,000 feet? 16,000 feet?
--------
Draw the picture.
You have a right triangle with base angle = 17 degrees
and side opposite = 10,000 ft.
-------
Find the hypotenuse:
sin(17) = 10,000/hypotenuse
------
hypotenuse = 10,000/sin(17) = 34,203.04 ft
-----
time = 34,203.04/345 = 99.14 seconds = 1 minute 39.14 seconds
----------------------
Note: Follow the same procedure for 16,000 ft.
---------------------------------------
Cheers,
Stan H.
==============